Velocity of a particle is expressed as
Vx=40-5t² m/s.
(a) find average acceleration.
(b)find instantaneous acceleration.
(C)Draw the graph of this question.
1
Expert's answer
2020-03-19T11:31:19-0400

(a) The average acceleration


"a_x=\frac{v_2-v_1}{t_2-t_1}=\frac{(40-5t_2^2)-(40-5t_1^2)}{t_2-t_1}""=-5(t_2+t_1)."

(b) The instantaneous acceleration


"a_x(t)=v_x'(t)=(40-5t^2)'=-10t."

(c)